1. How do you take a graph and generate an equation from it?
2. When comparing two linear equations, how can I tell if the corresponding lines are parallel? How can I tell if they are perpendicular?

Answer by khwang(438)

Given a system of two linear equations:
a1 x + b1 y = c1...(1),
a2 x + b2 y = c2...(2)
slope of (1): m1 = -a1/b1 (if b1 <> 0)
slope of (1): m2 = -a2/b2 (if b2 <> 0)

if a1/b1 = a2/b2 (same solpe) but is not equal to c1/c2 then the two lines
are parallel
if a1/b1 * a2/b2 = -1 (ie a1a2 = -b1b2 ) then the two lines
are perpendicular.
Try to think when b1 or b2 = 0